Given f(x) = 6x2 + 7x − 5 and g(x) = 4x − 8,find the following.
f(1) + g(1)

Answer by richard1234(7193)

f(1) + g(1) = (6*1^2 + 7*1 - 5) + (4*1 - 8) = 8 - 4 = 4.